# A Necessary Condition for the Submergence of Proto-Neutron Star Magnetic Fields by Supernova Fallback

### v1

@@ 3.2 Numerical setup
We set the radius of the outer boundary to $r_{\rm out}=10^{3}~{}{\rm km}$. The computational domain consists of $[r_{\rm PNS},r_{\rm out}]$. The number of numerical grid points is $N_{r}=16384$. The radius of the grid exponentially increases with $e^{x_{a}}$, where $x_{a}=\ln r_{\rm PNS}+a\times dX$, $a$ is the radial grid index, and $dX=(\ln r_{\rm out}-\ln r_{\rm PNS})/N_{r}$. The minimum grid width is $\Delta r\approx 280~{}{\rm cm}$, which is much smaller than the pressure scale height near the PNS surface $\sim 10^{5}{\rm cm}$ for $\rho\sim 10^{10}~{}{\rm g~{}cm^{3}}$ and $T\sim 10^{11}~{}{\rm K}$.
@@
[…]
We consider a magnetized and non-rotating PNS. The PNS initially has a magnetic field given by $B^{(\theta)}=B_{\rm dip}=B_{\rm PNS}(r/r_{\rm PNS})^{-3}$ and $B^{r}=B^{\phi}=0$. The parentheses in the indices of physical quantities denote the values in the static observer frame. This configuration models the fallback material accreting perpendicularly to the dipole magnetic field. We will discuss how the PNS rotation affects the submergence process in Section 5.1.
@@
In the ghost cells inside $r_{\rm PNS}$, we fix $B^{i}$ to the initial value and set $\rho$ and $p$ according to the zero-gradient boundary condition. We adopt the mirror boundary condition for $u^{r}$. The electrical conductivity inside the PNS is sufficiently high that the electric field can be considered negligible (Igoshev et al., 2021). Therefore, we set the numerical flux in the radial component of Equation (12) to 0 at $r=r_{\rm PNS}$. From the outer boundary at $r=r_{\rm out}$, we inject non-magnetized gas with a density of $\rho_{\rm out}$, a velocity of $v_{\rm ff}(r_{\rm out})$, and a temperature of $5\times 10^{7}~{}{\rm K}$ as the fallback material.
@@ 4.1 Time evolution of general structures
The left panel of Figure 3 presents the radial profiles of three different pressures in the fiducial model at $t=0.118~{}{\rm s}$: $p_{\rm mag}$ (blue), $p$ (gold), and $p_{\rm ram}=\rho(v^{(r)})^{2}$ (gray), where $v^{(r)}=u^{(r)}/u^{(t)}$. As mentioned in Figure 1, there are three distinct regions: the free-fall region, the post-shock region, and the PNS magnetosphere. These regions correspond to where $p_{\rm ram}$, $p$, and $p_{\rm mag}$ are the highest among the three, respectively. We define the magnetospheric radius $r_{\rm M}$ and the shock radius $r_{\rm sh}$ as the radii at which $p_{\rm mag}=p$ and $p=p_{\rm ram}$, respectively. We then obtain $r_{\rm M}=20~{}{\rm km}$ and $r_{\rm sh}=300~{}{\rm km}$ at this time (see dashed lines).
@@
Figure 4 displays time-sequenced images of the radial profiles of $p_{\rm ram}$ (left), $p$ (middle), and $p_{\rm mag}$ (right) in the fiducial model. For $t>t_{\rm shock}$, both $p_{\rm ram}$ and $p$ change discontinuously at $r=r_{\rm sh}$. In the early stage, $r_{\rm sh}$ expands over time. The dashed lines in the left panel indicate the times $t=t_{\rm shock}$ and $t=t_{\rm stall}$, where $t_{\rm stall}$ is the time when the expansion of $r_{\rm sh}$ stops (i.e., $v_{\rm sh}=0$). We write the ratio of $p_{\rm ram}+p$ to $p_{\rm mag}$ as $\beta$. The line in the right panel shows the radius where $\beta=1$. This radius decreases rapidly in the early time, indicating the compression of the PNS fields. The $\beta=1$ line lies in the vicinity of the PNS surface, showing the formation of the thin magnetosphere.
In the later stage ($t\geq t_{\rm stall}$), the accretion shock stalls and recedes due to the neutrino cooling. As the accreted mass accumulates above the PNS magnetosphere, the pressure just above it increases. According to the relation of $T=(3p/a_{\rm rad})^{1/4}$, $T$ at $r\sim r_{\rm M}$ also increases. Since neutrino cooling is more effective at a higher temperature, $t_{\rm cool}$ can become smaller than $t_{\rm dyn}$ at $r\sim r_{\rm M}$. The effective neutrino cooling reduces $p$ at $r\sim r_{\rm M}$, and the radiation pressure is no longer able to support gravity. For this reason, the post-shock region collapses. We will provide the dependence of the shock stalling time scale on $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$ in Section 4.3.
@@
In the non-magnetized cases, the post-shock region is formed in $r_{\rm PNS}<r<r_{\rm sh}$, while the free-fall region is located in $r>r_{\rm sh}$. The cooling time scale is shorter than the dynamical time scale for $r_{\rm PNS}\lesssim r\lesssim r_{\rm PNS}+H$ at $t=t_{\rm stall}$, which is consistent with the assumption in Chevalier (1989) (see Section 2).
@@ 4.2 Magnetospheric and shock radii
[…]
@@
Substituting $\dot{q}_{\rm pair}$ and $\dot{q}_{\rm URCA}$ into Equation (17), we obtain the following shock radii, respectively (c.f., Equation 8 and 9):
@@ 4.3 Shock stalling time scale, $t^{\prime}_{\rm stall}$
As described in Section 4.1, the density just above the magnetosphere begins to increase rapidly at $t=t_{\rm stall}$ due to neutrino cooling. This abrupt increase can potentially lead to the formation of the new crust that confines the PNS magnetosphere. As a measure of the waiting time for the new crust formation, we define the shock stalling time scale, $t^{\prime}_{\rm stall}=t_{\rm stall}-t_{\rm shock}$, and investigate its parameter dependence. Additionally, we aim to obtain a fitting formula of $t^{\prime}_{\rm stall}$ to discuss a necessary condition for the submergence of the PNS field.
@@ 4.5 Necessary condition for the submergence of the PNS magnetic field
We consider that the submergence of the PNS fields requires the abrupt increase in the gas density in the post-shock region. The abrupt increase occurs when a sufficient amount of accreted matter accumulates at $r\sim r_{\rm M}$ and the gas density and temperature become high enough to induce rapid cooling (that is why the shock starts to stall at that time; see Section 4.1). This may not be true when the supernova fallback ceases before the cooling becomes effective. As we have ignored the duration of the supernova fallback in the above analysis, we investigate this effect.
@@
Figure 6 displays time-sequenced images of radial profiles of the three different pressures. Due to the absence of the free-fall region, the post-shock gas expands to reduce its density and temperature (see the middle panel). The simulation does not show the abrupt increase in the density, suggesting that the PNS magnetosphere will not be submerged for $t^{\prime}_{\rm stall}>t_{\rm fb}$. Therefore, the condition $t^{\prime}_{\rm stall}<t_{\rm fb}$ needs to be satisfied for the submergence.
@@ 5.1 Implications for the Diversity in Young isolated NS
We propose possible solutions for the supernova fallback based on the following two conditions and discuss the diversity of young isolated NSs. The first condition is a necessary condition for the submergence of the PNS field (Section 4.5). The second condition is related to an effect of the PNS rotation, which is ignored in our numerical models. The rotating PNS can drive relativistic outflows and repel the fallback matter (e.g., Zhong et al., 2021). We take this effect into account based on the theoretical study of Zhong et al. (2021).
@@
The condition related to the PNS rotation is as follows (Zhong et al., 2021). The spin-down luminosity is expressed as (Spitkovsky, 2006)
@@
where $\Omega_{\rm PNS}=2\pi/P$ is the angular velocity of the PNS, $P$ is the spin period, and $\chi$ is the offset angle between the rotation and dipole axes. We take $\chi=0$ for simplicity. Zhong et al. (2021) assumed that the relativistic outflows and fallback matter typically encounters at the fallback radius, $r_{\rm fb}\sim(M_{\rm PNS}t_{\rm fb}^{2})^{1/3}$. From the condition of $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}(r_{\rm fb})$, we obtain (see also Shigeyama & Kashiyama, 2018, for more generalized case)
@@
Since the spin-down time scale, $\sim 10~{}{\rm yr}$, is much longer than $t_{\rm fb}$, the spin-down luminosity $L_{\rm spin}$ is regarded as a constant with time. We employ $P=0.02~{}{\rm s}$ as a fiducial value in Equation (27), which is shorter than the pulse periods observed in CCOs ($\sim 0.1~{}{\rm s}$, Enoto et al., 2019; Borghese & Esposito, 2023).
@@
Figure 7: Possible solutions of the supernova fallback illustrated as a function of fallback mass and time. We employ $B_{\rm PNS}=10^{13}~{}{\rm G}$ and $P=0.02~{}{\rm s}$. We show two cases of $r_{\rm PNS}=10~{}{\rm km}$ (left) and $20~{}{\rm km}$ (right). The thick and thin lines stand for $t^{\prime}_{\rm stall}=t_{\rm fb}$ (Equation 24) and $L_{\rm spin}=\dot{M}_{\rm fb}v^{2}_{\rm ff}$ (Equation 27), respectively. The fallback material submerges the PNS magnetosphere for $t^{\prime}_{\rm stall}<t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (red-shaded region). The fallback material is repelled by the spin-down power for $L_{\rm spin}>\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (blue-shaded region). The fallback material disturbs the PNS magnetosphere for $t^{\prime}_{\rm stall}>t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (orange-shaded region). We represent the critical point branching into the three different solutions as a black point. We consider that CCOs form in the red-shaded region, whereas the other types of NSs are born in the orange- and/or blue-shaded regions.
@@
Figure 7 displays the possible solutions for $B_{\rm PNS}=10^{13}~{}{\rm G}$. Here, we have used the initial value of the accretion rate for the fallback matter, $\dot{M}_{\rm fb}=(2/3)(M_{\rm fb}/t_{\rm fb})$ (see Equation 1). The thick and thin lines represent $t^{\prime}_{\rm stall}=t_{\rm fb}$ and $L_{\rm spin}=\dot{M}_{\rm fb}v^{2}_{\rm ff}$, respectively. In the region where $t^{\prime}_{\rm stall}<t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (red-shaded region), the supernova fallback can form a new crust and submerge the PNS magnetosphere. In the region where $L_{\rm spin}>\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (blue-shaded region), the fallback material is repelled by the dipole spin-down power and therefore does not interact with the PNS magnetosphere. In this case, the NS magnetic field is determined solely by internal amplifications during the core-collapse and PNS phases. In the region where $t^{\prime}_{\rm stall}>t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (orange-shaded region), the fallback accretion is halted before the PNS magnetosphere is submerged, allowing the magnetosphere to survive. During the fallback phase, the accreting material interacts with the PNS and disturbs its magnetosphere. We speculate that the disturbance leads to an external amplification of the magnetic field. We will discuss the amplification mechanisms induced by the supernova fallback later.
@@
Figure 7 shows that the critical points, which branch into three different solutions, are located around $(t_{\rm fb},~{}M_{\rm fb})\approx(2\times 10^{3}~{}{\rm s},~{}2\times 10^{-6}~{}{\rm M}_{\odot})$ for $r_{\rm PNS}=10~{}{\rm km}$ and $(t_{\rm fb},~{}M_{\rm fb})\approx(2\times 10^{3}~{}{\rm s},~{}10^{-4}~{}{\rm M}_{\odot})$ for $r_{\rm PNS}=20~{}{\rm km}$ (see black points). As $r_{\rm PNS}$ becomes large, a higher $M_{\rm fb}$ is required for the submergence. Although not shown in this figure, the red-shaded region becomes narrower for a stronger $B_{\rm PNS}$ within the plotted parameter range.
@@
The solutions in Figure 7 will provide insight into the diversity of the young isolated NSs. The CCOs can form in the red-shaded region, while the other types of the NSs are born in the orange- and/or blue-shaded regions. The solutions mean that when $B_{\rm PNS}=10^{13}~{}{\rm G}$ and $P=0.02~{}{\rm s}$, the NS population can be dominated by the CCOs within the typical ranges of $M_{\rm fb}=10^{-(4-1)}~{}{\rm M}_{\odot}$ and $t_{\rm fb}=10^{0-3}~{}{\rm s}$. This may suggest that the CCOs are common in the young isolated NSs. In fact, despite the small number of the known CCOs, observational studies implied that the birth rate of the CCOs is as large as the other types of the NSs (Gaensler et al., 2000).
We note that multidimensional effects, many of which are ignored in our simulations, will modify the solutions in Figure 7. Recent multidimensional simulations of the supernova explosion suggested that the fallback gas may have a large angular momentum (see Janka et al., 2022, and references therein). When such material accretes onto the PNS, an accretion disk will form around the PNS magnetosphere (Takahashi & Ohsuga, 2017; Inoue et al., 2023, 2024). The differential rotation of the disk opens up the PNS dipole field, which enhances $L_{\rm spin}$ (Parfrey & Tchekhovskoy, 2017). Furthermore, MHD instabilities are also important. The surface at $r_{\rm M}$, where the magnetic pressure balances the radiation pressure, can be unstable to the magnetic Rayleigh-Taylor instability (see e.g., Stone & Gardiner, 2007). This instability leads to dynamic mixing around $r\sim r_{\rm M}$, suppressing the submergence of the PNS magnetosphere (Bernal et al., 2013). Therefore, the solutions require further updates. However, we consider that our results will offer valuable guidance for investigating the diversity of young isolated NSs through multidimensional simulations.
@@
The supernova fallback may lead to an external amplification of the PNS magnetic field. When the fallback matter reaches the PNS without submerging the surface magnetic fields (orange-shaded region in Figure 7), it will chaotically disturb the magnetosphere. This chaotic process can lead to stronger and compressed multipolar magnetic fields (see discussion in Zhong et al., 2021). Furthermore, Barrère et al. (2022) proposed a magnetar formation scenario through the supernova fallback, in which the PNS field is amplified due to the Tayler-Spruit dynamo triggered by the PNS spin-up from the supernova fallback (see also Barrère et al., 2025). Although a detailed investigation of magnetic field evolution is beyond the scope of this study, we speculate that supernova fallback serves as an additional factor contributing to magnetic field amplification.

### v2

@@ 3.2 Numerical setup
We set the radius of the outer boundary to $r_{\rm out}=10^{3}~{\rm km}$. The computational domain consists of $r_{\rm PNS},r_{\rm out}]$. The number of numerical grid points is $N_{r}=16384$. The radius of the grid exponentially increases with $e^{x_{a}}$, where $x_{a}=\ln r_{\rm PNS}+a\times dX$, $a$ is the radial grid index, and $dX=(\ln r_{\rm out}-\ln r_{\rm PNS})/N_{r}$. The minimum grid width is $\Delta r\approx 280~{\rm cm}$, which is much smaller than the pressure scale height near the PNS surface $\sim 10^{5}{\rm cm}$ for $\rho\sim 10^{10}~{\rm g~cm^{3}}$ and $T\sim 10^{11}~{\rm K}$.
@@
We adopt typical values for the PNS mass and radius. The PNS contracts and cools down by neutrino emission after its birth. This contraction phase is considered to last for $\sim 1-10~{\rm s}$. The mass and radius of the PNS typically lie in $\approx 1-2~M_{\odot}$ and $\approx 10$$20~{\rm km}$, respectively (e.g., Fischer2010; Nagakura2022). We take $M_{\rm PNS}=1.4~M_{\odot}$ as a fiducial value. We adopt $r_{\rm PNS}=10$ and $20~{\rm km}$ and keep it fixed in the simulations for simplicity. We first show the results for $r_{\rm PNS}=10~{\rm km}$. We present the results for $r_{\rm PNS}=20~{\rm km}$ in Section 4.4 and explain how the results depend on $r_{\rm PNS}$.
We consider a magnetized and non-rotating PNS. The PNS initially has a magnetic field given by $B^{(\theta)}=B_{\rm init}=B_{\rm PNS}(r/r_{\rm PNS})^{-3}$ and $B^{r}=B^{\phi}=0$. The parentheses in the indices of physical quantities denote the values in the static observer frame. This configuration models the fallback material accreting perpendicularly to the dipole magnetic field. We ignore the effects of the large-scale field geometry and magnetic tension. We will discuss these effects in Section 5.2. We will also discuss how the PNS rotation affects the submergence process in Section 5.1.
@@
In the ghost cells inside $r_{\rm PNS}$, we fix $B^{i}$ to the initial value and set $\rho$ and $p$ according to the zero-gradient boundary condition.  [^1] The zero-gradient boundary condition at the inner boundary results in a discontinuity with the hydrostatic gas. To investigate the effect of the discontinuity, we conducted a non-accreting test. We confirm that the discontinuity does not affect the dynamics of the supernova fallback. Therefore, the inner boundary condition adopted in this study will not affect our conclusion. We adopt the mirror boundary condition for $u^{r}$. The electrical conductivity inside the PNS is sufficiently high that the electric field can be considered negligible (Igoshev2021). Therefore, we set the numerical flux in the radial component of Equation (12) to 0 at $r=r_{\rm PNS}$. From the outer boundary at $r=r_{\rm out}$, we inject non-magnetized gas with a density of $\rho_{\rm out}$, a velocity of $v_{\rm ff}(r_{\rm out})$, and a temperature of $5\times 10^{7}~{\rm K}$ as the fallback material.
@@ 4.1 Time evolution of general structures
The left panel of Figure 3 presents the radial profiles of three different pressures in the fiducial model at $t=0.118~{\rm s}$: $p_{\rm mag}$ (blue), $p$ (gold), and $p_{\rm ram}=\rho(v^{(r)})^{2}$ (gray), where $v^{(r)}=u^{(r)}/u^{(t)}$. As mentioned in Figure 1, there are three distinct regions: the free-fall region, the post-shock region, and the PNS magnetosphere. These regions correspond to where $_{\rm ram}, pp$, and $p_{\rm mag}$ are the highest among the three, respectively. We define the magnetospheric radius $r_{\rm M}$ and the shock radius $r_{\rm sh}$ as the radii at which $p_{\rm mag}=p$ and $p=p_{\rm ram}$, respectively. We then obtain $r_{\rm M}=20~{\rm km}$ and $r_{\rm sh}=300~{\rm km}$ at this time (see dashed lines).
@@
Figure 4 displays time-sequenced images of the radial profiles of $p_{\rm ram}$ (left), $p$ (middle), and $p_{\rm mag}$ (right) in the fiducial model. For $t>t_{\rm shock}$, both $p_{\rm ram}$ and $p$ change discontinuously at $r=r_{\rm sh}$. The dashed lines in the left panel indicate the times $t=t_{\rm shock}$ and $t=t_{\rm stall}$, where $t_{\rm stall}$ is the time when the expansion of $r_{\rm sh}$ stops (i.e., $v_{\rm sh}=0$). In the early stage ($t_{\rm shock}<t<t_{\rm stall}$), $r_{\rm sh}$ expands over time.
In the later stage ($t\geq t_{\rm stall}$), the accretion shock stalls and recedes due to the neutrino cooling. As the accreted mass accumulates above the PNS magnetosphere, the pressure just above it increases. According to the relation of $(3p/a_{\rm rad})^{1/4}, TT$ at $r\sim r_{\rm M}$ also increases. Since neutrino cooling is more effective at a higher temperature, $t_{\rm cool}$ can become smaller than $t_{\rm dyn}$ at $r\sim r_{\rm M}$. The effective neutrino cooling reduces $p$ at $r\sim r_{\rm M}$, and the radiation pressure is no longer able to support gravity. For this reason, the post-shock region collapses. We will provide the dependence of the shock stalling time scale on $\dot{M}_{\rm fb}$ and $B_{\rm PNS}$ in Section 4.3.
The line in the right panel of Figure 4 shows the radius of $\beta=1$, where $\beta=(p_{\rm ram}+p)/p_{\rm mag}$. This radius decreases rapidly in the early time, indicating the compression of the PNS fields. The $\beta=1$ line lies in the vicinity of the PNS surface for $t>0.05~{\rm s}$, showing the formation of a thin magnetosphere.
@@
In the non-magnetized cases, the post-shock region is formed in $r_{\rm PNS}<r<r_{\rm sh}$, while the free-fall region is located in $r>r_{\rm sh}$. The cooling time scale is shorter than the dynamical time scale for $r_{\rm PNS}\lesssim r\lesssim r_{\rm PNS}+H$ at $t=t_{\rm stall}$, which is consistent with the assumption in Chevalier1989 (see Section 2).
@@ 4.2 Magnetospheric and shock radii
We semi-analytically derive steady solutions of $r_{\rm M}$ and $r_{\rm sh}$. For simplicity, we use the spatial average of the magnetic field strength within the magnetosphere, denoted by $\bar{B}$. From the magnetic flux conservation law $\int^{\infty}_{r_{\rm PNS}}B_{\rm init}2\pi rdr=\int^{r_{\rm M}}_{r_{\rm PNS}}\bar{B}2\pi rdr$, we then obtain the magnetic pressure at $r=r_{\rm M}$:
@@
Substituting $\dot{q}_{\rm pair}$ and $\dot{q}_{\rm URCA}$ into Equation (17), we obtain the following shock radii, respectively (c.f., Equations 8 and 9):
@@ 4.3 Shock stalling time scale, $t^{\prime}_{\rm stall}$
As described in Section 4.1, the density just above the magnetosphere begins to increase rapidly at $t=t_{\rm stall}$ due to neutrino cooling. This abrupt increase can potentially lead to the formation of the new crust that confines the PNS magnetosphere. As a measure of the waiting time for the new crust formation, we define the shock stalling time scale, $t^{\prime}_{\rm stall}=t_{\rm stall}-t_{\rm shock}$, and investigate its parameter dependence. Additionally, we aim to obtain a fitting formula of $t^{\prime}_{\rm stall}$ to derive a necessary condition for the submergence of the PNS field.
@@ 4.5 Necessary condition for the submergence of the PNS magnetic field
We consider that the submergence of the PNS fields requires the abrupt increase in the gas density in the post-shock region. The abrupt increase occurs when a sufficient amount of accreted matter accumulates at $r\sim r_{\rm M}$, and the gas density and temperature become high enough to induce rapid cooling (that is why the shock starts to stall at that time; see Section 4.1). This may not be true when the supernova fallback ceases before the cooling becomes effective. As we have ignored the duration of the supernova fallback in the above analysis, we investigate this effect.
@@
Figure 6 displays time-sequenced images of radial profiles of the three different pressures. Due to the absence of the free-fall region, the accreted material redistributes itself to maintain the hydrostatic equilibrium. As a result, the post-shock gas expands to reduce its density and temperature (see the middle panel). The simulation does not show the abrupt increase in the density, suggesting that the PNS magnetosphere will not be submerged for $t^{\prime}_{\rm stall}>t_{\rm fb}$. Therefore, the condition $t^{\prime}_{\rm stall}<t_{\rm fb}$ needs to be satisfied for the submergence. The subsequent evolution of the hydrostatic atmosphere may become important for the submergence that occurs at a stage much later than the PNS phase. We will discuss the late-time submergence in Section 5.2.
@@ 5.1 Implications for the Diversity in Young isolated NS
We propose possible solutions for the supernova fallback based on the following two conditions and discuss the diversity of young isolated NSs. The first condition is a necessary condition for the submergence of the PNS field (Section 4.5). The second condition is related to an effect of the PNS rotation, which is ignored in our numerical models. The rotating PNS can drive relativistic outflows and repel the fallback matter (e.g., Zhong2021). We take this effect into account based on the theoretical study of Zhong2021.
@@
The condition related to the PNS rotation is as follows (Zhong2021). The spin-down luminosity is expressed as (Spitkovsky2006)
@@
where $\Omega_{\rm PNS}=2\pi/P$ is the angular velocity of the PNS, $P$ is the spin period, and $\chi$ is the offset angle between the rotation and dipole axes. We take $\chi=0$ for simplicity. Zhong2021 assumed that the relativistic outflows and fallback matter typically encounters at the fallback radius, $r_{\rm fb}\sim(M_{\rm PNS}t_{\rm fb}^{2})^{1/3}$. From the condition of $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}(r_{\rm fb})$, we obtain (see also Shigeyama2018, for more generalized case)
@@
Since the spin-down time scale, $\sim 10~{\rm yr}$, is much longer than $t_{\rm fb}$, the spin-down luminosity $L_{\rm spin}$ is regarded as a constant with time. We employ $P=0.02~{\rm s}$ as a fiducial value in Equation (27), which is shorter than the pulse periods observed in CCOs ($\sim 0.1~{\rm s}$, Enoto2019; Borghese2023).
@@
Figure 7: Possible solutions of the supernova fallback illustrated as a function of fallback mass and time. We employ $P=0.02~{\rm s}$. We show three cases of $(B_{\rm PNS},r_{\rm PNS})=(10^{13}~{\rm G},10~{\rm km})$ (left), $(B_{\rm PNS},r_{\rm PNS})=(10^{14}~{\rm G},10~{\rm km})$ (middle), and $(B_{\rm PNS},r_{\rm PNS})=(10^{13}~{\rm G},20~{\rm km})$ (right). The thick and thin lines stand for $t^{\prime}_{\rm stall}=t_{\rm fb}$ (Equation 24) and $L_{\rm spin}=\dot{M}_{\rm fb}v^{2}_{\rm ff}$ (Equation 27), respectively. The fallback material can submerge the PNS magnetosphere for $t^{\prime}_{\rm stall}<t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (red-shaded region). The fallback material is repelled by the spin-down power for $L_{\rm spin}>\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (blue-shaded region). The fallback material disturbs the PNS magnetosphere for $t^{\prime}_{\rm stall}>t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (orange-shaded region). We represent the critical point branching into the three different solutions as a black point. We consider that CCOs form in the red-shaded region, whereas the other types of NSs are born in the orange- and/or blue-shaded regions.
@@
Figure 7 displays the possible solutions. We have used the initial value of the accretion rate for the fallback matter, $\dot{M}_{\rm fb}=(2/3)(M_{\rm fb}/t_{\rm fb})$ (see Equation 1). Three cases are presented: $(B_{\rm PNS},r_{\rm PNS})=(10^{13}~{\rm G},10~{\rm km})$ (left), $(B_{\rm PNS},r_{\rm PNS})=(10^{14}~{\rm G},10~{\rm km})$ (middle), and $(B_{\rm PNS},r_{\rm PNS})=(10^{13}~{\rm G},20~{\rm km})$ (right). The thick and thin lines represent $t^{\prime}_{\rm stall}=t_{\rm fb}$ and $L_{\rm spin}=\dot{M}_{\rm fb}v^{2}_{\rm ff}$, respectively. In the region where $t^{\prime}_{\rm stall}<t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (red-shaded region), the supernova fallback can form a new crust and submerge the PNS magnetosphere. In the region where $L_{\rm spin}>\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (blue-shaded region), the fallback material is repelled by the dipole spin-down power and therefore does not interact with the PNS magnetosphere. In this case, the NS magnetic field is determined solely by internal amplifications during the core-collapse and PNS phases. In the region where $t^{\prime}_{\rm stall}>t_{\rm fb}$ and $L_{\rm spin}<\dot{M}_{\rm fb}v_{\rm ff}^{2}$ (orange-shaded region), the fallback accretion is halted before the PNS magnetosphere is submerged, allowing the magnetosphere to survive. During the fallback phase, the accreting material interacts with the PNS and disturbs its magnetosphere. We speculate that the disturbance leads to an external amplification of the magnetic field. We will discuss the amplification mechanisms induced by the supernova fallback later.
@@
Figure 7 shows that the critical points (black points), which branch into three different solutions, are located around $(t_{\rm fb},M_{\rm fb})\approx(10^{3}~{\rm s},10^{-6}~{\rm M}_{\odot})$ in the left panel, $(t_{\rm fb},M_{\rm fb})\approx(3\times 10^{2}~{\rm s},10^{-5}~{\rm M}_{\odot})$ in the middle panel, and $(t_{\rm fb},M_{\rm fb})\approx(10^{3}~{\rm s},7\times 10^{-5}~{\rm M}_{\odot})$ in the right panel. From the left and middle panels, the red-shaded region becomes narrower for a stronger $B_{\rm PNS}$ within the plotted parameter range. In addition, the $M_{\rm fb}$ required for the submergence increases as $r_{\rm PNS}$ increases (see left and right panels).
@@
The solutions in Figure 7 will provide insight into the diversity of the young isolated NSs. The CCOs can form in the red-shaded region, while the other types of the NSs are born in the orange- and/or blue-shaded regions. The solutions mean that when $B_{\rm PNS}=10^{13}~{\rm G}$ and $P=0.02~{\rm s}$, the NS population can be dominated by the CCOs within the typical ranges of $M_{\rm fb}=10^{-(4-1)}~{\rm M}_{\odot}$ and $t_{\rm fb}=10^{0-3}~{\rm s}$. This may suggest that the CCOs are common in the young isolated NSs (deLuca2008; Torres2016). In fact, despite the small number of the known CCOs, observational studies implied that the birth rate of the CCOs is as large as the other types of the NSs (Gaensler2000).
Figure 7 allows us to assess whether the necessary condition derived in this study is consistent with previous studies. Here, we refer to Ho2011 and Vigano2012 for comparison. Their approach differs from ours in that they focused on the long-term evolution of crustal magnetic fields (i.e., submerged fields), rather than on neutrino cooling in the fallback flows. They solved the induction equation and obtained the re-emerged magnetic field strength at the NS surface as a function of time. By comparing their result with CCO observations, they inferred a fallback mass of $M_{\rm fb}\gtrsim 10^{-3}~{\rm M}_{\odot}$ for $(B_{\rm PNS},r_{\rm PNS})\approx(10^{14}~{\rm G},10~{\rm km})$. The middle panel of Figure 7 shows that their inferred value exceeds $M_{\rm fb}$ on the $t^{\prime}_{\rm stall}=t_{\rm fb}$ line. Thus, the necessary condition is not contradict with their models.
@@
The supernova fallback may lead to an external amplification of the PNS magnetic field. When the fallback matter reaches the PNS without submerging the surface magnetic fields (orange-shaded region in Figure 7), it will chaotically disturb the magnetosphere. This chaotic process can lead to stronger and compressed multipolar magnetic fields (see discussion in Zhong2021). Furthermore, Barrere2022 proposed a magnetar formation scenario through the supernova fallback, in which the PNS field is amplified due to the Tayler-Spruit dynamo triggered by the PNS spin-up from the supernova fallback (see also Barrere2025). Although a detailed investigation of magnetic field evolution is beyond the scope of this study, we speculate that supernova fallback serves as an additional factor contributing to magnetic field amplification.